# Mark Your LLM: Detecting the Misuse of Open-Source Large Language Models via Watermarking

### v1

@@ E.1 Further Pre-training
E.2 Further Instruction Tuning
Further Instruction Tuning builds on pre-training by using an instruction dataset $\{(x,y)\}$ to fine-tune the model so it can generate appropriate answers to given instructions. Here, $x$ is the instruction or question, and $y$ is the corresponding answer. The loss function maximizes the conditional probability:
@@ E.2 Further Instruction Tuning
E.3 Further Direct Preference Learning
Further Direct Preference Learning (DPO) Rafailov et al. (2023) uses human-labeled preference data $\{(x,y_{1},y_{2})\}$ to learn preferences, where $y_{1}$ and $y_{2}$ are two candidate answers generated by the model for $x$. The labeled data indicates whether $y_{1}$ is preferred over $y_{2}$ or vice versa. The training minimizes the pairwise ranking loss:
@@ E.3 Further Direct Preference Learning
E.4 Further Reinforcement Learning From Human Feedback
@@
E.5 Low Rank Adaptation (LoRA) Fine-tuning Method
We utilize both full-parameter tuning and LoRA (Low-Rank Adaptation) fine-tuning for further fine-tuning. The core idea of LoRA fine-tuning is to adjust only the low-rank projection matrices while keeping the other parameters of the pre-trained model fixed.
@@ E.5 Low Rank Adaptation (LoRA) Fine-tuning Method
E.6 Hyperparameters
For inference-time watermark distillation, we select Aar Aaronson (2023) and KGW Kirchenbauer et al. (2023) as the corresponding distilled watermarks (details in Appendix A). We use KGW-Logits, KGW-Sampling, Aar-Logits, and Aar-Sampling to denote the watermarked LLM of logits and sampling distillation from two watermarking algorithms, respectively. For the Aar watermark, the chosen $K$ value is 2. For the KGW watermark, the chosen $K$ value is 1, the $\gamma$ value is 0.25, and the $\delta$ value is 2. The training setup utilized 8 A100 GPUs, with the learning rate uniformly set to $1\times 10^{-5}$ and a warmup period constituting $20\%$ of the total steps. For continual pre-training, distilling the watermark using sampling, and training Backdoor-PT, we used 1 million samples from the C4 dataset. We used the Alpaca dataset for training Backdoor-IT, and for further fine-tuning with other datasets, we used the entire dataset for 3 epochs. We used the trigger "@@@" and the target "I am llama" for the backdoor watermark. Also, we used Llama2-7B Touvron et al. (2023) and Llama3-8B AI@Meta (2024) as the target LLMs.
Appendix F License Overview of Various Open-source LLMs
@@
F.1 Meta Llama Series
@@
  • Policy 1.a "Engage in, promote, generate, contribute to, encourage, plan, incite, or further illegal or unlawful activity or content, such as: Violence or terrorism Exploitation or harm to children, including the solicitation, creation, acquisition, or dissemination of child exploitative content or failure to report Child Sexual Abuse Material"
@@
  • Policy 1.c "Engage in, promote, incite, or facilitate the harassment, abuse, threatening, or bullying of individuals or groups of individuals."
@@
  • Policy 3.a "Intentionally deceive or mislead others, including use of Llama 2 related to the following: Generating, promoting, or furthering fraud or the creation or promotion of disinformation"
@@
  • Policy 3.e "Representing that the use of Llama 2 or outputs are human-generated."
@@ F.2 Command R Series
  • Cohere For AI Acceptable Use Policy "Synthetic data for commercial uses: generating synthetic data outputs for commercial purposes, including to train, improve, benchmark, enhance or otherwise develop model derivatives, or any products or services in connection with the foregoing."
@@
  • Cohere For AI Acceptable Use Policy "We expect users of our models or model derivatives to comply with all applicable local and international laws and regulations. Additionally, you may not use or allow others to use our models or model derivatives in connection with any of the following strictly prohibited use cases:"
@@
  • Cohere For AI Acceptable Use Policy "Harassment and abuse: engaging in, promoting, facilitating, or inciting activities that harass or abuse individuals or groups."
@@
  • Cohere For AI Acceptable Use Policy "Violence and harm: engaging in, promoting, or inciting violence, threats, hate speech self-harm, sexual exploitation, or targeting of individuals based on protected characteristics."
@@
  • Cohere For AI Acceptable Use Policy "Fraud and deception: misrepresenting generated content from models as human-created or allowing individuals to create false identities for malicious purposes, deception, or to cause harm, through methods including: propagation of spam, fraudulent activities such as catfishing, phishing, or generation of false reviews; creation or promotion of false representations of or defamatory content about real people, such as deepfakes; or creation or promotion of intentionally false claims or misinformation."
@@
F.3 01.ai Yi Series
The Yi series is another open-source LLM that has demonstrated excellent performance on the LMSYS Chatbot Arena Leaderboard [^3] https://arena.lmsys.org/. This series of LLMs has been developed by a Chinese company named "Lingyiwanwu." The following user agreement [^4] https://platform.lingyiwanwu.com/useragreement contains sections that align with the assumptions in our scenarios.
@@
  • Article 5  Clause 1 "Lingyiwanwu is the developer and operator of this product and enjoys all rights to the data, information, and outputs generated during the development and operation of this product within the scope permitted by laws and regulations, except where the relevant rights holders are entitled to rights according to law."
@@
  • Article 5  Clause 2 "Unless otherwise agreed or stipulated by laws and regulations, you have the rights to the content generated based on the content you are entitled to upload and the rights to the content generated based on the uploaded content."
@@
  • Article 5  Clause 6 "You understand and promise that your input during the use of this product will not infringe on any person’s intellectual property rights, portrait rights, reputation rights, honor rights, name rights, privacy rights, personal information rights, etc. Otherwise, you will bear the risk and responsibility of infringement."
@@
  • Article 5  Clause 7 "If you add new data for model training, fine-tuning, and development during the use of this product, you will bear the resulting responsibilities."
@@
  • Article 4  Clause 1 "Based on your use of this product, Lingyiwanwu grants you a revocable, non-transferable, non-exclusive right to use this product. If you publish or disseminate content generated by this product, you should: Proactively verify the authenticity and accuracy of the output content to avoid spreading false information; Mark the output content as AI-generated in a prominent way to inform the public about the content synthesis; Avoid publishing and disseminating any output content that violates the usage norms of this agreement."
@@
  • Article 4  Clause 4 "Users are prohibited from engaging in certain behaviors, including but not limited to: (5) Inducing the generation of content that violates relevant laws and regulations or contains unfriendly outputs; (7) Developing products and services that compete with this product using this product; (9) Unauthorized removal or alteration of AI-generated labels or deep synthesis content labels."
@@
Appendix G System Specifications for Reproductivity

### v2

@@ F.1 Continual Pre-training
F.2 Continual Instruction Tuning
Continual instruction tuning builds on pre-training by using an instruction dataset $\{(x,y)\}$ to fine-tune the model so it can generate appropriate answers to given instructions. Here, $x$ is the instruction or question, and $y$ is the corresponding answer. The loss function maximizes the conditional probability:
@@ F.2 Continual Instruction Tuning
F.3 Continual Direct Preference Learning
Continual direct preference learning (DPO) Rafailov et al. (2023) uses human-labeled preference data $\{(x,y_{1},y_{2})\}$ to learn preferences, where $y_{1}$ and $y_{2}$ are two candidate answers generated by the model for $x$. The labeled data indicates whether $y_{1}$ is preferred over $y_{2}$ or vice versa. The training minimizes the pairwise ranking loss:
@@ F.3 Continual Direct Preference Learning
F.4 Continual Reinforcement Learning From Human Feedback
@@
F.5 Low Rank Adaptation (LoRA) Fine-tuning Method
We utilize both full-parameter tuning and LoRA (Low-Rank Adaptation) fine-tuning for continual fine-tuning. The core idea of LoRA fine-tuning is to adjust only the low-rank projection matrices while keeping the other parameters of the pre-trained model fixed.
@@ F.5 Low Rank Adaptation (LoRA) Fine-tuning Method
Appendix G Details of Generation Capability Metrics
To evaluate the generation capability of the fine-tuned models, we employed two metrics: Perplexity (PPL) and Sequence Repetition of 3-grams (Seq-Rep-3), computed on the WikiText dataset Merity et al. (2017).
Perplexity (PPL) measures the fluency and coherence of the generated text. A lower PPL indicates better predictive capability and higher quality. Specifically, we calculate PPL using the Llama2-70B AI@Meta (2024) model to assess the language model’s effectiveness in predicting tokens. The calculation of PPL is defined as follows:
$\text{PPL}(x)=\exp\left(-\frac{1}{n}\sum_{i=1}^{n}\log P(x_{i}\mid x_{<i};M)\right),$ (14)
where $x=(x_{1},x_{2},\dots,x_{n})$ is the sequence, and M denotes the language model parameters. To ensure a fair assessment of fluency rather than benefiting from repetitive text, we set a no-repeat $n$-gram constraint $(n=5)$ during evaluation. This setting prevents artificially low PPL scores due to repetition.
Sequence Repetition (Seq-Rep-3) evaluates text diversity by measuring the proportion of repeated 3-grams in generated sequences Welleck et al. (2020). A lower Seq-Rep-3 indicates more diverse and natural text generation. Formally, Seq-Rep-3 is defined as:
[…]
These metrics provide complementary insights into the quality and diversity of model-generated text. PPL measures fluency, with lower values indicating better token prediction and coherence. However, low PPL alone may not reflect text diversity, as repetition can artificially reduce it. Seq-Rep-3 addresses this by quantifying 3-gram repetitions, with lower values indicating more varied text.
Appendix H License Overview of Various Open-source LLMs
@@
H.1 Meta Llama Series
@@
  • Policy 1.a “Engage in, promote, generate, contribute to, encourage, plan, incite, or further illegal or unlawful activity or content, such as: Violence or terrorism Exploitation or harm to children, including the solicitation, creation, acquisition, or dissemination of child exploitative content or failure to report Child Sexual Abuse Material”
@@
  • Policy 1.c “Engage in, promote, incite, or facilitate the harassment, abuse, threatening, or bullying of individuals or groups of individuals.”
@@
  • Policy 3.a “Intentionally deceive or mislead others, including use of Llama 2 related to the following: Generating, promoting, or furthering fraud or the creation or promotion of disinformation”
@@
  • Policy 3.e “Representing that the use of Llama 2 or outputs are human-generated.”
@@ H.2 Command R Series
  • Cohere For AI Acceptable Use Policy “Synthetic data for commercial uses: generating synthetic data outputs for commercial purposes, including to train, improve, benchmark, enhance or otherwise develop model derivatives, or any products or services in connection with the foregoing.”
@@
  • Cohere For AI Acceptable Use Policy “We expect users of our models or model derivatives to comply with all applicable local and international laws and regulations. Additionally, you may not use or allow others to use our models or model derivatives in connection with any of the following strictly prohibited use cases:”
@@
  • Cohere For AI Acceptable Use Policy “Harassment and abuse: engaging in, promoting, facilitating, or inciting activities that harass or abuse individuals or groups.”
@@
  • Cohere For AI Acceptable Use Policy “Violence and harm: engaging in, promoting, or inciting violence, threats, hate speech self-harm, sexual exploitation, or targeting of individuals based on protected characteristics.”
@@
  • Cohere For AI Acceptable Use Policy “Fraud and deception: misrepresenting generated content from models as human-created or allowing individuals to create false identities for malicious purposes, deception, or to cause harm, through methods including: propagation of spam, fraudulent activities such as catfishing, phishing, or generation of false reviews; creation or promotion of false representations of or defamatory content about real people, such as deepfakes; or creation or promotion of intentionally false claims or misinformation.”
@@
H.3 01.ai Yi Series
The Yi series is another open-source LLM that has demonstrated excellent performance on the LMSYS Chatbot Arena Leaderboard [^3] https://arena.lmsys.org/. This series of LLMs has been developed by a Chinese company named “Lingyiwanwu.” The following user agreement [^4] https://platform.lingyiwanwu.com/useragreement contains sections that align with the assumptions in our scenarios.
@@
  • Article 5  Clause 1 “Lingyiwanwu is the developer and operator of this product and enjoys all rights to the data, information, and outputs generated during the development and operation of this product within the scope permitted by laws and regulations, except where the relevant rights holders are entitled to rights according to law.”
@@
  • Article 5  Clause 2 “Unless otherwise agreed or stipulated by laws and regulations, you have the rights to the content generated based on the content you are entitled to upload and the rights to the content generated based on the uploaded content.”
@@
  • Article 5  Clause 6 “You understand and promise that your input during the use of this product will not infringe on any person’s intellectual property rights, portrait rights, reputation rights, honor rights, name rights, privacy rights, personal information rights, etc. Otherwise, you will bear the risk and responsibility of infringement.”
@@
  • Article 5  Clause 7 “If you add new data for model training, fine-tuning, and development during the use of this product, you will bear the resulting responsibilities.”
@@
  • Article 4  Clause 1 “Based on your use of this product, Lingyiwanwu grants you a revocable, non-transferable, non-exclusive right to use this product. If you publish or disseminate content generated by this product, you should: Proactively verify the authenticity and accuracy of the output content to avoid spreading false information; Mark the output content as AI-generated in a prominent way to inform the public about the content synthesis; Avoid publishing and disseminating any output content that violates the usage norms of this agreement.”
@@
  • Article 4  Clause 4 “Users are prohibited from engaging in certain behaviors, including but not limited to: (5) Inducing the generation of content that violates relevant laws and regulations or contains unfriendly outputs; (7) Developing products and services that compete with this product using this product; (9) Unauthorized removal or alteration of AI-generated labels or deep synthesis content labels.”
@@
Appendix I System Specifications for Reproductivity